# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning

### v3

@@ 5 Analysis of Language alignment and Effects on multilingualism
From layers 18–28 Merlin reaches its peak sentence–level retrieval@5 (~0.30). The rise coincides with the PEFT fine–tuning in Stage II, which adjusts a small set of low-rank weights on top of the frozen LLM while keeping the connector $f_{\sigma^{\star}}$ fixed. This adaptation strengthens the multilingual abstractions that naturally emerge in the decoder’s middle blocks—the region reported to be the most language-agnostic once parallel supervision is available (liu2025middlelayerrepresentationalignmentcrosslingual). As a result, non-English queries are already embedded in the English reasoning sub-space before the later, generation-oriented layers are reached, providing a solid cross-lingual scaffold for downstream mathematical reasoning tasks such as AfriMGSM.
@@ 6 Related Work
LRLs often suffer from limited annotated data and high-quality machine translation resources, creating significant challenges for NLP models (magueresse2020lowresourcelanguagesreviewpast; lee_pre-trained_2022; app13158566; deshpande2024chainoftranslationpromptingcotrnovel; qin2024multilinguallargelanguagemodel; tonja-etal-2024-nlp). Moreover, these languages can exhibit complex morphological structures and orthographic variations (ghosh2024morphologybasedinvestigationpositionalencodings; nzeyimana2024lowresourceneuralmachinetranslation; lopo2024constructingexpandinglowresourceunderrepresented; issaka2024ghanaiannlplandscapelook; lusito2023textnormalizationlowresourcelanguages), posing further challenges. While multilingual models can handle many languages, they frequently underperform in truly low-resource settings and may even trail behind strong monolingual baselines (yoon-etal-2024-langbridge; du2024chinesetinyllmpretraining; blevins-etal-2024-breaking). Existing efforts to boost LRL performance focus primarily on dataset creation (adelani-etal-2021-masakhaner; muhammad2022naijasentinigeriantwittersentiment; adelani-etal-2024-sib; yong2024lexcgengeneratingdataextremely) or adaptation of pretrained models (alabi-etal-2022-adapting; wu2024adaptinglargelanguagemodels; csaki2023efficientlyadaptingpretrainedlanguage; schmidt-etal-2024-self), yet these strategies remain constrained by data scarcity and fail to fully address complex linguistic properties (urbizu-etal-2023-scaling; wang2023challengingbenchmarklowresourcelearning).
@@
Recent efforts to enhance multilingual reasoning in LLMs include carefully devised prompt engineering to facilitate multi-step inference or chain-of-thought reasoning (huang-etal-2023-languages; qin-etal-2023-cross), although these techniques often lose effectiveness with smaller open-source models (touvron2023llama2openfoundation). Another approach involves supervised fine-tuning on translated data (chen2024breakinglanguagebarriersmultilingual; zhu2024questiontranslationtrainingbetter), which infuses language or task-specific knowledge into the model. Building upon this, fan2025slamefficientmultilingualreasoning extend the work of zhu2024questiontranslationtrainingbetter by optimizing both the training parameters and the training stages focusing on language-specific neurons. Their improvements lead to higher average performance on multilingual mathematical reasoning benchmarks. However, reliance on machine translation can introduce errors (shi2022languagemodelsmultilingualchainofthought) and may overlook the LLM’s built-in multilingual capabilities. To circumvent these issues, MindMerger (huang2024mindmergerefficientboostingllm) sidesteps autoregressive translation by directly integrating an external multilingual encoder, preserving the LLM’s internal parameters.
@@ 7 Conclusion
Our analysis shows that model merging enhances cross-lingual alignment across layers compared to using LLMs alone. In particular, our framework, Merlin, improves alignment within the model’s internal reasoning subspace, yielding stronger performance on reasoning tasks. We also find that the choice of encoder is critical for low-resource languages: encoders with greater exposure to LRL data produce representations that align more closely with the LLM’s internal language (often English), thereby supplying richer cross-lingual signals for model merging.
@@ Appendix D Ablation of Training Adapters
Table 8 contrasts two parameter-efficient adapters applied to the decoder. LoRA (hu2022lora) injects trainable low-rank update matrices into the linear weights, while all original parameters stay frozen. DoRA (liu2024doraweightdecomposedlowrankadaptation) keeps the same low-rank structure but factorises each weight into a direction and a magnitude, yielding the same footprint yet a more explicit decomposition.
@@
Table 8: Ablation of LoRA vs. DoRA adapters for Merlin (Gemma-2 9 B + NLLB-600 M). Mrl. = mean over mid-resource languages (Bn, Th, Sw); Hrl. = mean over seven high-resource languages; averages are arithmetic means of the shown scores.
@@ Appendix E AmericasNLI Evaluation
To assess whether the mapping curriculum generalises beyond the African and European test beds, we evaluate Merlin on AmericasNLI, a three-way NLI benchmark that targets Aymara, Guarani, and Quechua. For Merlin and the MindMerger baseline, we follow the same data pipeline as for AfriXNLI: 3,000 English premise–hypothesis pairs are sampled from MultiNLI, translated into each target language with NLLB 600m distilled, and used as supervision for Stage I.b, Stage I.c, and Stage II. [^3] No task-specific data from the original AmericasNLI development set is seen during training. The non-adapted LLaMAX2-7B-XNLI (lu-etal-2024-llamax) serves as the baseline decoder-only model. Results are summarised in Table 9.
@@ Appendix F Training Configurations
Gemma-2 backbone <bos><start_of_turn>user Below is an instruction that describes a task. Write a response that appropriately completes the request. {query} Let’s think step by step. <end_of_turn><start_of_turn>model
@@
Both Merlin and MindMerger are trained on identical corpora. Stage I.a uses 9,000 generic English–target sentence pairs per language drawn from the NLLB-200 corpus. Stage I.b is supplied with up to 3,000 automatically translated questions per language, also generated with NLLB. For Stage I.c and the subsequent Task-Specialisation stage, task-specific data are employed: the MGSM track relies on the translated MetaMathQA split, whereas the AfriXNLI track is trained with translated MultiNLI.

### v4

@@ 5 Analysis of Language Alignment and Effects on Multilingualism
From layers 18–28 Merlin reaches its peak sentence–level retrieval@5 (~0.30). The rise coincides with the PEFT fine–tuning in Stage II, which adjusts a small set of low-rank weights on top of the frozen LLM while keeping the connector $f_{\sigma^{\star}}$ fixed. This adaptation strengthens the multilingual abstractions that naturally emerge in the decoder’s middle blocks—the region reported to be the most language-agnostic once parallel supervision is available (Liu and Niehues, 2025). As a result, non-English queries are already embedded in the English reasoning sub-space before the later, generation-oriented layers are reached, providing a solid cross-lingual scaffold for downstream mathematical reasoning tasks such as AfriMGSM.
Isolating the Effects of Curriculum Structure and Data Volume
To disentangle curriculum effects from data volume, we compare Merlin and MindMerger under an identical supervision budget of 12k examples per language. As shown in Table 5, Merlin still consistently outperforms MindMerger on both MGSM (78.3 vs. 73.4) and MSVAMP (80.4 vs. 78.3), despite using no additional data.
Since both methods use the same total supervision, the gains cannot be attributed to data quantity. Instead, the key difference lies in how supervision is organized: MindMerger concentrates most data in a single bilingual mapping stage, whereas Merlin distributes supervision across multiple curriculum stages that progressively transition from general alignment to task-specific reasoning. The consistent improvements across mid- and high-resource languages indicate that this staged curriculum yields more effective cross-lingual and task-aligned representations, validating curriculum structure as the primary source of Merlin’s gains.
Model Bn Th Sw Ja Zh De Fr Ru Es En Mrl. Hrl. Avg
MGSM
MERLIN (1k / stage) 76.4 77.6 74.4 71.6 76.0 78.4 77.6 81.6 83.6 85.6 76.1 79.2 78.3
MSVAMP
MERLIN (1k / stage) 73.0 75.7 79.4 79.7 82.2 82.9 82.7 81.6 82.8 84.4 76.0 82.3 80.4
Table 5: Full per-language comparison between MindMerger and Merlin using the same total amount of supervision (12k pairs) per language on MGSM and MSVAMP. MindMerger uses 9k bilingual pairs (mapping) + 3k QA (task-aware), whereas Merlin (1k / stage) uses 9k bilingual pairs + 1k for each of Stage Ib, Stage Ic, and Stage II. Bn/Th/Sw are mid-resource (MRL); Ja/Zh/De/Fr/Ru/Es/En are high-resource (HRL).
Disentanglement between Understanding and Reasoning abilities of LLM
Since MGSM and AfriMGSM are direct translations of GSM8K, all models generate the final answer in English. This means the underlying reasoning task is identical across languages, so any performance gap cannot stem from limitations in the LLM’s reasoning ability. Instead, it directly reflects the quality of cross-lingual alignment. The above results strongly support this interpretation: English accuracy remains stable across all methods, while Merlin delivers the largest gains in low- and mid-resource languages. Merlin also improves substantially on AfriXNLI, a setting focused on sentence-level understanding rather than multi-step reasoning, further indicating that the curriculum strengthens cross-lingual representations rather than altering the LLM’s reasoning behavior. Our layer-wise retrieval and encoder-choice analyses reinforce this conclusion by showing clearer alignment in the decoder’s intermediate layers.
@@ 6 Related Work
LRLs often suffer from limited annotated data and high-quality machine translation resources, creating significant challenges for NLP models (Magueresse et al., 2020; Lee et al., 2022; Khan et al., 2023; Deshpande et al., 2024; Qin et al., 2024; Tonja et al., 2024). Moreover, these languages can exhibit complex morphological structures and orthographic variations (Ghosh et al., 2024; Nzeyimana, 2024; Lopo and Tanone, 2024; Issaka et al., 2024; Lusito et al., 2023), posing further challenges. While multilingual models can handle many languages, they frequently underperform in truly low-resource settings and may even trail behind strong monolingual baselines (Yoon et al., 2024; Du et al., 2024; Blevins et al., 2024). Existing efforts to boost LRL performance focus primarily on dataset creation (Adelani et al., 2021; Muhammad et al., 2022; Adelani et al., 2024a; Yong et al., 2024) or adaptation of pretrained models (Alabi et al., 2022; Wu et al., 2024; Csaki et al., 2023; Schmidt et al., 2024), yet these strategies remain constrained by data scarcity and fail to fully address complex linguistic properties (Urbizu et al., 2023; Wang et al., 2023).
@@
Recent efforts to enhance multilingual reasoning in LLMs include carefully devised prompt engineering to facilitate multi-step inference or chain-of-thought reasoning (Huang et al., 2023; Qin et al., 2023), although these techniques often lose effectiveness with smaller open-source models (Touvron et al., 2023). Another approach involves supervised fine-tuning on translated data (Chen et al., 2024; Zhu et al., 2024), which infuses language or task-specific knowledge into the model. Building upon this, Fan et al. (2025) extend the work of Zhu et al. (2024) by optimizing both the training parameters and the training stages focusing on language-specific neurons. Their improvements lead to higher average performance on multilingual mathematical reasoning benchmarks. However, reliance on machine translation can introduce errors (Shi et al., 2022) and may overlook the LLM’s built-in multilingual capabilities. To circumvent these issues, MindMerger (Huang et al., 2024) sidesteps autoregressive translation by directly integrating an external multilingual encoder, preserving the LLM’s internal parameters.
@@ 7 Conclusion
Our analysis shows that model merging enhances cross-lingual alignment across layers compared to using LLMs alone. In particular, our framework, Merlin, improves alignment within the model’s internal reasoning subspace, yielding stronger performance on reasoning tasks even under tightly matched supervision budgets. We also find that the choice of encoder is critical for low-resource languages: encoders with greater exposure to LRL data produce representations that align more closely with the LLM’s internal language (often English), thereby supplying richer cross-lingual signals for model merging.
@@ 8 Limitations
Acknowledgments
This research was supported in part by the Natural Sciences and Engineering Research Council (NSERC) of Canada. David Adelani acknowledges the funding of IVADO and the Canada First Research Excellence Fund. We acknowledge the use of generative AI tools for improving proofreading and readability, no scientific content was generated by the tools.
@@ Appendix D Ablation of Training Adapters
Table 9 contrasts two parameter-efficient adapters applied to the decoder. LoRA (Hu et al., 2022) injects trainable low-rank update matrices into the linear weights, while all original parameters stay frozen. DoRA (Liu et al., 2024) keeps the same low-rank structure but factorises each weight into a direction and a magnitude, yielding the same footprint yet a more explicit decomposition.
@@
Table 9: Ablation of LoRA vs. DoRA adapters for Merlin (Gemma 2 9 B + NLLB-600 M). Mrl. = mean over mid-resource languages (Bn, Th, Sw); Hrl. = mean over seven high-resource languages; averages are arithmetic means of the shown scores.
@@ Appendix E AmericasNLI Evaluation
To assess whether the mapping curriculum generalises beyond the African and European test beds, we evaluate Merlin on AmericasNLI, a three-way NLI benchmark that targets Aymara, Guarani, and Quechua. For Merlin and the MindMerger baseline, we follow the same data pipeline as for AfriXNLI: 3,000 English premise–hypothesis pairs are sampled from MultiNLI, translated into each target language with NLLB 600m distilled, and used as supervision for Stage I.b, Stage I.c, and Stage II. [^4] No task-specific data from the original AmericasNLI development set is seen during training. The non-adapted LLaMAX2-7B-XNLI (Lu et al., 2024) serves as the baseline decoder-only model. Results are summarised in Table 10.
@@ Appendix F Training Configurations
Gemma 2 backbone <bos><start_of_turn>user Below is an instruction that describes a task. Write a response that appropriately completes the request. {query} Let’s think step by step. <end_of_turn><start_of_turn>model
@@
Both Merlin and MindMerger are trained on identical corpora. Stage I.a uses 9,000 generic English–target sentence pairs per language drawn from the NLLB-200 corpus. Stage I.b is supplied with up to 3,000 automatically translated questions per language, also generated with NLLB. For Stage I.c and the subsequent Task-Specialization stage, task-specific data are employed: the MGSM track relies on the translated MetaMathQA split, whereas the AfriXNLI track is trained with translated MultiNLI.